OpenStudy (anonymous):

fill in the missing term so that the quadratic equation has a graph that opens up with a vertex of (- 3, -25), and x intercepts at x= -8 and x=2. ( do not include the negative sign in your answer.)

OpenStudy (mertsj):

\[y=ax^2+bx+c\] 0=a(2)^2+b(2)+c 0=a(-8)^2+b(-8)+c -25=a(-3)^2+b(-3)+c

OpenStudy (mertsj):

Solve the system

OpenStudy (mertsj):

\[4a^2+2b+c\]=0 \[64a-8b+c=0\] \[9a-3b+c=-25\]